# Spectrum Projection Score: Aligning Retrieved Summaries with Reader Models in Retrieval-Augmented Generation

### v1

@@ 5.2 Main Experiment
Table 2 summarises the effectiveness of our proposed xCompress framework across five QA datasets and two retrieval-based generation paradigms (text-to-text and text-to-embedding). Results demonstrate that incorporating our Spectrum Projection Score (SPS) consistently improves downstream Exact Match (EM) and F1 scores over the baseline compression methods. For instance, on the NQ dataset using the LLAMA 3.1 model, SPS improves performance from 35.2/47.49 (EM/F1) to 39.4/51.18, highlighting SPS’s effectiveness in selecting retrieval summaries that align better with the reader model’s internal representations. Our framework consistently outperforms perplexity-based (PPL and LongPPL) methods, except in two cases involving the Qwen3 model on the HotpotQA and NQ datasets. In these instances, selecting summaries based on PPL yielded slightly better performance than SPS. We attribute this to the Qwen model’s overconfidence on these datasets, possibly due to substantial overlap between its pretraining data and these evaluation datasets, a hypothesis supported by recent literature suggesting dataset contamination or repeated exposure during training (wu2025reasoning). Further, empirical analysis revealed significantly lower entropy, approximately three times lower, in Qwen’s predictions on HotpotQA and NQ compared to LLAMA 3.1, reinforcing that Qwen likely recalls answers directly from memorised content  [^4] Experiment result details are in Appendix C.. Consequently, summaries chosen via perplexity reflect familiar, memorised content rather than optimal alignment, paradoxically enhancing performance but reducing generalisability.
@@ 6.2 How Spectrum Projection Score Performs with Different Sentence Embeddings?
We analyse how the number of generated summaries influences the effectiveness of our proposed Spectrum Projection Score (SPS). As shown in Figure 5, xCompress surpasses the Compact baseline (yoon-etal-2024-compact) (EM 62.4, F1 71.25) by generating just two summaries, highlighting immediate gains from summary diversity. Performance improves notably up to five summaries (EM 65.4, F1 73.11), beyond which additional summaries yield marginal benefits. Thus, we select five summaries as the optimal balance between performance and computational efficiency.
@@ 7 Conclusion
We propose the Spectrum Projection Score (SPS), a training-free, representation-level metric that evaluates the semantic alignment between retrieved summaries and the reader model’s internal geometry. Building on SPS, we introduce xCompress, an inference-time controller that guides summary selection through test-time sampling and adaptive filtering. Extensive experiments across five QA benchmarks and multiple LLMs demonstrate that SPS outperforms perplexity-based baselines in both correlation with answer quality and downstream task performance.
@@ B.3 Further Discussion
To analyse this further, we adopt the commonly used assumption that token embeddings follow a multi-dimensional Gaussian distribution (ng2011sparse; karvonen2025saebench). Under this assumption, we consider the ratio between the mean vector and the bounded vector, defined as:

### v2

@@ 5.2 Main Experiment
Table 2 summarises the effectiveness of our proposed xCompress framework across five QA datasets and two retrieval-based generation paradigms (text-to-text and text-to-embedding). Results demonstrate that incorporating our Spectrum Projection Score (SPS) consistently improves downstream Exact Match (EM) and F1 scores over the baseline compression methods. For instance, on the NQ dataset using the LLAMA 3.1 model, SPS improves performance from 35.2/47.49 (EM/F1) to 39.4/51.18, highlighting SPS’s effectiveness in selecting retrieval summaries that align better with the reader model’s internal representations. Our framework consistently outperforms perplexity-based (PPL and LongPPL) methods, except in two cases involving the Qwen3 model on the HotpotQA and NQ datasets. In these instances, selecting summaries based on PPL yielded slightly better performance than SPS. We attribute this to the Qwen model’s overconfidence on these datasets, possibly due to substantial overlap between its pretraining data and these evaluation datasets, a hypothesis supported by recent literature suggesting dataset contamination or repeated exposure during training (Wu et al. 2025). Further, empirical analysis revealed significantly lower entropy, approximately three times lower, in Qwen’s predictions on HotpotQA and NQ compared to LLAMA 3.1, reinforcing that Qwen likely recalls answers directly from memorised content (Experiment details are in the Appendix C.). Consequently, summaries chosen via perplexity reflect familiar, memorised content rather than optimal alignment, paradoxically enhancing performance but reducing generalisability.
@@ 6.2 How Spectrum Projection Score Performs with Different Sentence Embeddings?
We analyse how the number of generated summaries influences the effectiveness of our proposed Spectrum Projection Score (SPS). As shown in Figure 5, xCompress surpasses the Compact baseline (Yoon et al. 2024) (EM 62.4, F1 71.25) by generating just two summaries, highlighting immediate gains from summary diversity. Performance improves notably up to five summaries (EM 65.4, F1 73.11), beyond which additional summaries yield marginal benefits. Thus, we select five summaries as the optimal balance between performance and computational efficiency.
@@ 7 Conclusion
We propose the Spectrum Projection Score (SPS), a training-free, representation-level metric that evaluates the semantic alignment between retrieved summaries and the reader model’s internal geometry, moving beyond perplexity and naive similarity as proxies for utility. Building on SPS, we introduce xCompress, an inference-time controller that guides summary selection through reader-guided test-time sampling, ranking, and adaptive filtering. Extensive experiments across five QA benchmarks and multiple LLMs demonstrate that SPS not only outperforms perplexity-based baselines in correlation with answer quality and downstream task performance, but also offers a principled, model-agnostic tool for diagnosing and improving retrieval–reader compatibility in RAG systems.
Acknowledgments
This work was supported by King’s Computational Research, Engineering, and Technology Environment (CREATE) and was supported in part by the UK Engineering and Physical Sciences Research Council (EPSRC) through a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2) and a New Horizons grant (grant no. EP/X019063/1), and KCL’s Impact Acceleration Account (grant no. EP/X525571/1). A PhD studentship from the Chinese Scholarship Council funds Zhanghao Hu.
@@ B.3 Further Discussion
To analyse this further, we adopt the commonly used assumption that token embeddings follow a multi-dimensional Gaussian distribution (Ng et al. 2011; Karvonen et al. 2025). Under this assumption, we consider the ratio between the mean vector and the bounded vector, defined as: